# GLINT-RU: Gated Lightweight Intelligent Recurrent Units for Sequential Recommender Systems

### v1

@@ 5. Related Works
Traditional Sequential Recommendation Models Transformers, RNNs are popular effective frameworks for sequential recommendation. Various studies have been conducted to leverage transformers for recommendation tasks. Transformer-based SASRec (Kang and McAuley, 2018) makes predictions based on multi-head attention within long-/short-term interactions. SSE-PT (Wu et al., 2020) improves the performance by adding user embeddings to the transformer. MB-STR (Yuan et al., 2022) is also a variant of transformer-based model, which captures diverse user behavior dynamics, and alleviates the impacts from data sparsity problems. RNN-based GRU4Rec (Hidasi et al., 2015) learns dependencies among the items using gated recurrent units. HRNN (Quadrana et al., 2017) as an advanced RNN-based model adds an additional GRU layer to learn information from user sessions and instantly track the user preferences.
Efficient Recommendation Models To tackle the high computational complexity of effective transformer-based SRSs, and accelerate the inference speed for real-world applications, researchers endeavor to invent increasingly effcient models. AutoSeqRec (Liu et al., 2023b) is constructed based on Autoencoder, which is a innovative work that captures long-term preferences through collaborative filtering. To further improve the model efficiency, DMAN (Tan et al., 2021) combines the long-term attention net and recurrent attention net to memorize users’ interests dynamically and support efficient inference. LinRec (Liu et al., 2023a) cuts down the computational cost of transformer-based backbones by changing the dot product in the attention mechanism. LightSAN (Fan et al., 2021) projects the initial interactions into representations with shorter length, which is also an efficient approach for transformer-based models. SSM-based models like Mamba4Rec (Liu et al., 2024) and RecMamba (Yang et al., 2024) utilizes selective state space models to achieve high performances and high efficiency, becoming emerging powerful tools for sequential recommendation tasks. FMLP-Rec (Zhou et al., 2022) with learnable filters and SMLP4Rec (Gao et al., 2024) with diverse mixers are representative pure MLP-based efficient SRSs. LRURec (Yue et al., 2024) is constructed based on linear recurrent units, and achieves fast inference bu employing recursive parallelization.
@@ 6. Conclusion
In this paper, we have presented an innovative dense selective GRU framework GLINT-RU for sequential recommendation tasks. Due to the parallel netork design and the implementation of efficient dense selective GRU, the computational cost can be substantially reduced, thereby resulting in faster inference speed compared with state-of-the-art efficient models. Additionally, our model GLINT-RU tackles the issue of low accuracy of traditional GRU frameworks, and represents a significant advancement in the efficient recommendation models. GLINT-RU applies gate mechanisms to every module, deeply filtering and selecting the information according to the data environment. Dense GRU allows each hidden state to contain the information from the adjacent items, which captures local temporal features of user behavior to enhance the predictive accuracy. Our extensive experiments demonstrate that GLINT-RU achieves outstanding performance, not only improving the model accuracy, but also accelerating the training and inference speed dramatically. These results underscore our GLINT-RU’s potential to become a novel, stable, and efficient framework on datasets with various sparsity and sequence length. As an efficient and accurate recommender system that defeats the state-of-the-art efficient RNN-, transformer-, MLP- and SSM-based recommendation models, we believe that our novel framework will become a valuable foundation for research in more domains in recommender systems.

### v2

@@ 5. Related Works
Traditional Sequential Recommendation Models Transformers and RNNs are widely recognized as popular and effective frameworks for sequential recommendation tasks. Numerous studies have been conducted to explore the potential of transformers in enhancing recommendation performance. The transformer-based SASRec (Kang and McAuley, 2018) predicts user preferences by leveraging multi-head attention to model both long- and short-term interactions. SSE-PT (Wu et al., 2020) further improves recommendation accuracy by incorporating user embeddings into the transformer architecture. Similarly, MB-STR (Yuan et al., 2022), a transformer-based variant, captures diverse user behavior dynamics and effectively mitigates the challenges posed by data sparsity. On the other hand, RNN-based methods like GRU4Rec (Hidasi et al., 2015) model dependencies among items through gated recurrent units, enabling them to learn sequential patterns. HRNN (Quadrana et al., 2017), an advanced RNN-based framework, integrates an additional GRU layer to extract session-level information and dynamically track user preferences over time. In comparison to these models, GLINT-RU demonstrates superior capability in capturing high-quality semantic features and fine-grained positional representations, leading to improved model accuracy and performance.
Efficient Recommendation Models To tackle the high computational complexity of existing SRSs, and accelerate the inference speed for real-world applications, researchers endeavor to invent increasingly efficient models. AutoSeqRec (Liu et al., 2023b) is constructed based on Autoencoder, which is an innovative work that captures long-term preferences through collaborative filtering. To further improve the model efficiency, DMAN (Tan et al., 2021) combines the long-term attention net and recurrent attention net to memorize users’ interests dynamically and support efficient inference. LinRec (Liu et al., 2023a) cuts down the computational cost of transformer-based backbones by changing the dot product in the attention mechanism. SSM-based models like Mamba4Rec (Liu et al., 2024a) and RecMamba (Yang et al., 2024) utilize selective state space models to achieve high performances and high efficiency, becoming emerging powerful tools for sequential recommendation tasks. FMLP-Rec (Zhou et al., 2022) with learnable filters and SMLP4Rec (Gao et al., 2024) with diverse mixers are representative pure MLP-based efficient SRSs. LRURec (Yue et al., 2024) is constructed based on linear recurrent units and achieves fast inference by employing recursive parallelization. These models often require stacked network structures to achieve better results, and their performance is not stable enough when faced with different sequence lengths. However, GLINT-RU only needs one layer to achieve stable high performance.
@@ 6. Conclusion
In this paper, we have presented an innovative dense selective GRU framework, GLINT-RU for sequential recommendation tasks. Due to the paralleled network design and implementation of efficient dense selective GRU with linear complexity, the computational cost can be substantially reduced, resulting in state-of-the-art inference speed. Additionally, our GLINT-RU models improve the quality of semantic features and fine-grained positional information for recommendation tasks. Gate mechanisms are widely applied in GLINT-RU, deeply filtering and selecting information. It uses a dense selective GRU that aggregates information from adjacent items and generates high-quality latent item representations based on dependencies with fine-grained positional information to learn context information. By mixing dense selective GRU with linear attention, GLINT-RU can capture important interactions and item dependencies simultaneously. Our extensive experiments demonstrate that GLINT-RU achieves outstanding performance, not only improving model accuracy but also accelerating training and inference speed dramatically. These results underscore GLINT-RU’s potential to become a novel, stable, and efficient framework in various scenarios. As an efficient recommender system that defeats state-of-the-art models, we believe our novel framework will become a valuable foundation.
Acknoledgement
This research was partially supported by Research Impact Fund (No.R1015-23), APRC - CityU New Research Initiatives (No.9610565, Start-up Grant for New Faculty of CityU), CityU - HKIDS Early Career Research Grant (No.9360163), Hong Kong ITC Innovation and Technology Fund Midstream Research Programme for Universities Project (No.ITS/034/22MS), Hong Kong Environmental and Conservation Fund (No. 88/2022), and SIRG - CityU Strategic Interdisciplinary Research Grant (No.7020046), Collaborative Research Fund (No.C1043-24GF), Huawei (Huawei Innovation Research Program), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Ant Group (CCF-Ant Research Fund, Ant Group Research Fund), Alibaba (CCF-Alimama Tech Kangaroo Fund No. 2024002), CCF-BaiChuan-Ebtech Foundation Model Fund, and Kuaishou.
Appendix A Model Training Methods
In this section, we introduce the details of the training methods of GLINT-RU, including the loss function and the optimization algorithm of GLINT-RU displayed through pseudo-code.
Based on the prediction score in Eq.(13), we adopt the cross entropy loss function defined below for training GLINT-RU:
where are the ground-truth labels of the user-item interactions. Now we can train GLINT-RU by following the guidance in Algorithm 1. As can be seen in the optimization algorithm, the training process of GLINT-RU can be easily implemented. The paralleled expert linear attention and dense selective GRU can be balanced by updating mixing weights. Additionally, to help GLINT-RU find a better local optimal solution, we follow the suggestion of (Glorot and Bengio, 2010) to initialize the model parameters , which helps prevents the vanishing/exploding gradient problem.
Input: User-item interactions , corresponding ground-truth labels
Output: Well-trained weights of model
Appendix B Hyperparameter Settings
In this section, we will introduce the hyperparameter settings of our GLINT-RU and the baselines, and offer implementation details for reproducibility.
The parameter settings of all baselines are displayed in Table 5. In addition to the hyperparameter settings in Table 5, we also set the number of interests in LightSANs to 5, as the the average sequence length is less than 10, which is shown in Table 1. We set the expansion factor in Mamba4Rec as 2, and set the expansion factor in SMLP4Rec as 1 to avoid the GPU memory occupation error. The feature selection in SMLP4Rec is consistent with the settings in the original paper. (Gao et al., 2024). For ML-1M dataset, we choose [’genre’, ’movie title’, ’release year’] as the input features in the SMLP structure. And for Amazon Beauty and Games, we select [’categories’, ’brand’] as the input features. These features are employed to enhance the prediction performance from the item ids in SMLP4Rec.
| Method | Recall@10 | MRR@10 | NDCG@10 |
| Default (SiLU) | 0.4472 | 0.2498 | 0.2964 |
| ReLU | 0.4293 | 0.2300 | 0.2770 |
| Sigmoid | 0.4328 | 0.2319 | 0.2793 |
| ELU | 0.4308 | 0.2321 | 0.2789 |
Appendix C Ablation Study: Activation Functions
Unlike ReLU, which has a sharp kink at zero, SiLU is smooth and continuously differentiable. This smoothness facilitates better gradient flow during backpropagation, leading to faster and more stable training. Additionally, SiLU mitigates the vanishing gradient problem more effectively than Sigmoid and Tanh, as its gradient does not saturate as quickly, allowing for more efficient training of deeper networks. Incorporating SiLU as the activation function in the gate introduces non-linearity while retaining important item information and filtering out irrelevant item information, which is advantageous for complex tasks. Studies have demonstrated that SiLU enhances training stability and performance in deep learning models. In practical applications, user-item interactions often contain noise, as some selected items may be irrelevant. Under such conditions, employing SiLU typically achieves higher accuracy compared to traditional activation functions like ReLU and ELU. To further verify the effectiveness of the SiLU activation function, we conducted additional experiments where SiLU was replaced by ELU, Sigmoid, and ReLU in the dense selective gate. The results are presented in Table 6.
Appendix D Parameter Analysis: Layer Number
In this section, we conduct the parameter analysis on the number of layers to observe the model’s accuracy and efficiency in stacked GLINT-RU. We change the number of GLINT-RU layers from 1 to 4, and evaluate the model performance with NDCG@10 using the dataset Amazon Beauty. The results are shown in Figure 6.
From Figure 6, it is evident that as the number of layers increases, the accuracy of the GLINT-RU model exhibits a relatively gradual upward trend. This indicates that stacking GLINT-RU layers contributes positively to improving accuracy. However, this improvement is not pronounced, and the model’s performance may even slightly decline when the number of layers becomes excessively large. Additionally, we analyze the model efficiency of GLINT-RU in Table 7:
| Inference Time | Training Speed | GPU Memory | |
| 1 | 278ms | 3.8s/epoch | 2.62GB |
| 2 | 397ms | 6.2s/epoch | 5.17GB |
| 3 | 501ms | 8.9s/epoch | 6.97GB |
| 4 | 610ms | 11.8s/epoch | 8.86GB |
The results in Table 7 indicate that as the number of layers increases, the efficiency of GLINT-RU decreases rapidly. Combined with the previous model performance analysis, we can conclude that under resource constraints, using only one layer of efficient GLINT-RU can achieve high model efficiency and accuracy simultaneously.
Appendix E Ablation Study: Amazon Beauty Dataset and Video Games Dataset
To further verify the effectiveness of the components in GLINT-RU, we conducted more ablation studies on the Amazon Beauty and Amazon Video Games datasets. The results are shown in Tables 9 and 8. The results demonstrate that the gated GRU module plays a critical role in capturing dependencies among interactions and fine-grained positional representations. The linear attention mechanism enhances the model’s ability to capture interactions between relevant items within the sequence. Adding a temporal convolution layer improves performance by incorporating contextual information from adjacent items. Finally, the gated MLP block filters complex information from the expert mixing block, further contributing to the overall performance. These findings indicate that all components of GLINT-RU are effective and collectively contribute to its high performance on both the Amazon Beauty and Amazon Video Games datasets.
| Method | Recall@10 | MRR@10 | NDCG@10 |
| Default (GLINT-RU) | 0.6573 | 0.3549 | 0.4266 |
| w/o GRU | 0.6379 | 0.3125 | 0.4023 |
| w/o Linear Attention | 0.6459 | 0.3375 | 0.4106 |
| w/o Temporal Conv1d | 0.6443 | 0.3351 | 0.4084 |
| w/o Gated MLP | 0.6416 | 0.3339 | 0.4068 |
| Method | Recall@10 | MRR@10 | NDCG@10 |
| Default (GLINT-RU) | 0.4472 | 0.2498 | 0.2964 |
| w/o GRU | 0.4239 | 0.2298 | 0.2755 |
| w/o Linear Attention | 0.4296 | 0.2314 | 0.2781 |
| w/o Temporal Conv1d | 0.4246 | 0.2290 | 0.2751 |
| w/o Gated MLP | 0.4265 | 0.2308 | 0.2768 |